OpenStudy (this):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? (Round your answer to the nearest tenth of a second.) a) .6 seconds to 1.8 seconds b) .4 seconds to 1.9 seconds c) .2 seconds to 1.3 seconds d) .5 seconds to 2.1 seconds

OpenStudy (anonymous):

Don't you think we have to take derivative of the equation first.???

OpenStudy (eyust707):

no that would give you velocity

OpenStudy (shane_b):

It's already a position equation...

OpenStudy (eyust707):

with the quadratic you will obtain two solutions

OpenStudy (eyust707):

one time on the way up on on the way down

OpenStudy (eyust707):

the difference between the two is the solution

OpenStudy (eyust707):

nice

OpenStudy (shane_b):

I'm lazy

OpenStudy (anonymous):

the answer is c)

OpenStudy (eyust707):

I guess the ref is 5 ft tall

OpenStudy (shane_b):

must be the WNBA

OpenStudy (eyust707):

haha

OpenStudy (this):

there fore so the answer is C hun?

OpenStudy (eyust707):

yep

OpenStudy (shane_b):

Definitely C.

OpenStudy (campbell_st):

you have a question where you need to solve fo t then h(t) = 9 so you have 9 = -16t^2 + 24t + 5 or -16t^2 + 24t - 4 = 0 so solve for t lots of ways... but here is the long version using the general Quadratic formula \[t = \frac{-24 \pm \sqrt{24^2 - 4\times(-16\times(-4)}}{2\times(-16)}\] simplifying gives \[t = \frac{-24 \pm \sqrt{576 - 256}}{-32}\] after some further simplification you end up with \[t = \frac{-3 \pm \sqrt{5}}{-4}\] so you need to evaluate \[t = \frac{-3 - \sqrt{5}}{-4} .... and ... t=\frac{-3 + \sqrt{5}}{-4}\] its between these 2 time values that the ball is above 9 ft

OpenStudy (campbell_st):

t = 0.19 or rounding to 2 d.p t = 0.2 sec and t = 1.3 sec

OpenStudy (this):

thnks guys a lot ya all r the best :)

OpenStudy (callisto):

Same question... I got the same numerical answer as @campbell_st, but I don't know why I was wrong.. The arrangement of the choices was different. http://openstudy.com/users/callisto#/updates/4fdca110e4b0f2662fd27820

OpenStudy (shane_b):

I don't see why it's wrong either...it's just a simple quadratic equation and you can plainly see the answer on the graph.

OpenStudy (callisto):

Very true indeed!!